Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,100 helmets, using 2,201 kilograms of plastic. The plastic cost the company $16,728.

According to the standard cost card, each helmet should require 0.62 kilograms of plastic, at a cost of $8.00 per kilogram.

Required:

1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,100 helmets?

2. What is the standard materials cost allowed (SQ × SP) to make 3,100 helmets?

3. What is the materials spending variance?

4. What is the materials price variance and the materials quantity variance?

(For requirements 3 and 4, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations.)
Answer #1
1.

Each helmet requires 0.62 kilogram, So for 3,100 helmets
Standard quantity of kilograms of plastic = 3,100 X 0.62 = 1,922 kilograms
2.

Each helmet requires 0.62 kilogram X 8 = $ 4.96 per helmet, So for 3,100 helmets
Standard quantity of kilograms of plastic = 3,100 X 4.96 = $ 15,376
3.

Actual cost = $ 16,728
Standard = $ 15,376
Spending variance = 1,352 unfavorable
4.

Material price variance = (Actual Qty X Actual Price) - (Actual Qty X Standard Price)
                                                = 16,728 - (2,201 X 8) = (880) Favorable

Material quantity variance = (Actual Qty X Std Price) - (Std Qty X Standard Price)
                                                       = (2,201 X 8) - (1,922 X 8) = 2,232 Unfavorable